Suppose the domestic supply and demand curves for MP3 players in the United States are given by the following set of equations:

QS = 25 + 10P
QD = 925 – 5P

In the absence of international trade in MP3 players, how many MP3 players will be sold in the United States?
Answer #1

Answer
the equilibrium is at Qd=Qs in the case of absence of trade
equating both equations
25+10P=925-5P
15P=900
P=60
Q=25+10*60
Q=625
number of MP3 players will be sold in the United States is 625 units